# Discovery of Volatile Gas in the Giant Impact Disk around the 150-Myr old HD 23514

### v1

@@ A.2 Dust Emissivity & Silica Properties
To further validate our approach, we also generated dust emissivity functions for a young star called LkCa 15. The dust emissivity in LkCa 15 was classified as an ISM-like profile by Watson et al. (2009) where the emission is dominated by pristine silicate dust grains (i.e., amorphous and sub-micron). Therefore, we used its mid-IR archival spectra from Spitzer/IRS and JWST/MRS as the input and constructed the pseudocontinuum using the polynomial function for the former but a combination of two blackbodies for the latter. The resultant two dust emissivity functions agree well within a few percent in the 10 µm complex (5–14 µm), but in the 20 µm complex the JWST one is slightly higher by 20–40% likely due to the limited long-wavelength in determining the pseudocontinuum. The derived JWST LkCa 15 dust emissivity is shown in the right panel of Figure 1. Because our focus is in deriving the crystalline mass fraction indices by referencing the ISM-like profile in the 10 µm complex, the difference in the 20 µm region has no impact on the result. We used the JWST LkCa 15 emissivity function to compute the indices ( for pyroxene, for olivine and for silica) as defined by Watson et al. (2009). Note that these indices are defined as the normalized ratio in dust emissivity at a certain wavelength range covering the narrow features, the normalization of the reference (ISM-like) emissivity (i.e., the blue curve in the left panel of Figure 1) has no effects on the results. For HD 23514, the derived indices are: 2.54, 0.91, and 1.8, suggesting a high fraction of crystallinity preferentially in the pyroxene and silica mineralogy. We also derived the dust emissivity using the Spitzer IRS spectrum, and found similar indices within the uncertainty (corroborating no change in the dust composition).
@@
To quantitatively characterize the silica feature in HD 23514, we also derived the dust emissivity using the Spitzer IRS spectrum of T51 (a few Myr-old star). T51 is one of the five silica-rich PPDs characterized by Sargent et al. (2009) who found that the system lacks sub-µm amorphous dust and the crystalline silicate mass fraction is 15% including both pyroxene and olivine types, and the silica dust is best scribed as annealed silica although obsidian or amorphous quartz give equally good fits. Our derived indices for T51 are: 2.02, 1.50, and 2.36, respectively, consistent with the finding in HD 23514 above. Although the 10 µm feature in T51 and HD 23514 both peaks at 9.1 µm (Figure 1 right), the FWHM of the 10 µm feature in T51 is 1.5 times larger, suggesting the dominant size in HD 23514 is smaller, likely sub-µm.
@@
Silica is not abundant in the ISM. Its presence in circumstellar disks requires special conditions for formation like giant impacts as we discuss here. Space weathering, as discussed by Sargent et al. (2009), is one way to deposit silica by converting silicate minerals; however, it is unlikely to be the primary mechanism responsible for the large amount of sub-um size silica seen in HD 23514. Space weathering is a long-term process. µm-size silicate grains in the inner region of HD 23514 have short lifetimes (blowout size is 1 µm, and will be gone on orbital timescales), implying that converting small silicates, produced by collisional cascades, to silica is an inefficient process. For stars just emerge from the protoplanetary disk stage, massive flare events could speed up the process as suggested by Lisse et al. (2020), however, HD 23514 is not very active at an old age of 150 Myr (Appendix A.1). Comets could bring a large amount of fresh dust into the inner region during a period of high delivery rates like the late heavy bombardment, and create silica through space weathering. For solar system comets, the dust mass is 1014 g on average (Harker et al., 2023), meaning a total of 10 billion comets is needed to account for the dust observed in HD 23514 (1024 g derived in Appendix D). The most challenging problem in the comet delivery scenario is that the typical dust size is large, 100 µm (Rinaldi et al., 2017), i.e., space weathering should produce large-size silica, not sub-. In summary, the proposed giant impact scenario is the most likely explanation for the large amount of sub-µm silica grains in HD 23514.
Appendix B Other Mid-infrared Data and 3–5 µm Photometric Monitoring
@@
HD 23514 has various space- and ground-based mid-IR photometry, with the earliest measurement from IRAS in 1983 (Moshir et al., 1992; Ishihara et al., 2010; Rhee et al., 2008; Wright et al., 2010; Meng et al., 2012). Figure 7 summarizes these photometric observations between 5 and 25 µm. The plotted data have been color corrected using the Spitzer/IRS spectrum, assuming that it is a good representation of the SED shape in the given wavelength range over the period of interest (1983–2010). Because WISE 12 µm band was selected to match approximately that of IRAS (Cutri et al., 2012), they trace the same part of the 10 µm feature. The color corrected flux densities in these WISE and IRAS bands are 102.94.8 mJy and 131.618.4 mJy, respectively, i.e. the two fluxes agree within the 1.5 photometric uncertainty.
@@
IR variability within 20–30% levels, as an additional hallmark commonly found in the extreme debris disks (Moór et al., 2021), is known for HD 23514 (Meng et al., 2012). HD 23514 has been intensively monitoring by Spitzer/IRAC and WISE at 3–5 µm with the former in 3–day cadence when available, while the latter in 6–month cadence. Following the analysis method used for another IR variable, the RZ Psc disk (Su et al., 2023), Figures 8 and 9 show the short- and long-term disk fluxes at 3–5 µm between 2008 and 2024 where the two epochs of the IR spectra were also marked.
@@
Using these 3–5 µm data, the HD 23514 disk exhibits stochastic, both weekly and yearly IR variability with a weak, positive trend between the excess flux and color temperature. The inferred color temperatures are 700–800 K, consistent with the warmer dust temperature inferred from the featureless pseudocontinuum in Appendix A.2. The typical peak-to-peak flux variation (Figure 8) is 15% in both the 3.6 and 4.5 µm bands, more than 10 significance compared to the photometry stability (1%) determined by the field stars in the data. The 14-year long WISE data (Figure 9) show a similar behavior with the largest peak-to-peak flux difference of 25 mJy in the W2 band. Taking the typical (10 mJy) and maximum (25 mJy) 4.5 µm flux change at a face value, the minimal corresponding area change in dust cross section is (1–2.5) au2 by adopting a dust temperature of 800 K and a distance of 139 pc under optically thin assumption. It is interesting to note that this minimal area change in the dust cross section is similar to the hot CO2 gas emitting area (Table 1). The cross section changes can also be translated to the minimal change in the dust mass (assuming static state collisional cascades down to the radiation blowout size of 1 µm). The very lower limit on the associated dust mass is on the order of 3–7 g, corresponding to the total disruption of a 120–165 km-size asteroid (assuming a density of 3.5 g cm-3). This mass/size is a lower limit because of our assumptions (optically thin and static state).
Appendix C Gas model and Other Related Calculation
C.1 LTE Slab Model
@@
To fit the line emission, we adopted the slab-modeling Python package iris (Munoz-Romero et al., 2023) assuming the gas is in local LTE, with line data from the HITRAN database (Gordon et al., 2022). The line width was estimated with thermal broadening based on the temperature of the gas. In iris, the fitting is performed with the Bayesian nested sampling Python package dynesty (Speagle, 2020). We set the stopping criterion by the change in the remaining evidence (marginal likelihood ) when , where = with as the number of parameters in the model.
We restricted the fitting within the wavelength range of 14.0–17.0 µm for the emission of CO2, and adopted uniform priors for emitting area () and column density (), with and . The prior for the gas temperature () was set to be uniform initially, and later set as a normal distribution centered around 500 K with a standard deviation of around 200 K to speed up the search. We included 13CO2 with CO2 in our fitting, and assumed the emissions were from the same region ( and are the same) and the 12C/13C ratio was the same as ISM value (, e.g., Milam et al., 2005). The posterior distributions for the CO2 fitting is shown in Figure 10. The 15.4 µm 13CO2 line is the only (i.e., strongest) isotopologue line present in the data. Excluding it produces a slightly poor fit at 15.4 µm as shown in Figure 2. The best fit CO2 model parameters were summarized in Table 1 by assuming two different turbulence velocities: 0 km s-1 as the nominal and 2 km s-1 as the upper limit value. The two cases make very little difference in the resulting fits visually, illustrating the associated uncertainties in the MCMC fits might be underestimated, and reflecting the limitations of using the simple LTE slab model.
@@
Although there are hints of other molecules (CO, H2O and NH3) present in the continuum subtracted spectra (Figure 12), we did not attempt to determine their precise parameters using the slab models due to high noise and weak lines. Instead, we assumed they share similar model parameters as CO2, and adjust (and , if necessary) to achieve reasonable fits visually. The mid-infrared emission from molecules is due to ro-vibrational or vibrational transitions and is sensitive to their temperatures, i.e., the lines get significantly weaker as the temperature gets cooler. It is then meaningless to explore lower temperatures (or other parameters) without high S/N data. Keeping the same temperature for both H2O and NH3, the column density needs to be reduced by a factor of 2.2 and 1.4, respectively, in order to produce similar flux levels of the lines between 6.5 and 9.5 µm. The low abundance in both H2O and NH3 is consistent with their very short photodissociation lifetimes (Figure 5), by 2–3 order of magnitudes compared to that of CO2, without shielding. The model parameters for CO, which has the longest photodissociation lifetimes near HD 23514, are very uncertain due to the limited wavelength coverage from MIRI/MRS. The slab model has great difficulty to reproduce the observed line width if CO has a similar as CO2 (termed low case in the top panel of Figure 11). Alternatively, the broad width of CO lines can roughly be reproduced by an extra broadening of 200 km s-1 and increasing to 3000 K (termed high case in the middle panel of Figure 11). Such a high temperature can be reached at a stellocentric distance of 0.015 au around HD 23514 where its Keplerian velocity () is 277 km s-1. This seems to suggest the additional broadening might be related to the Keplerian motion of the CO gas. If the gas is indeed in a disk form, resolving the double-peak profiles in the CO emission could help to constrain the inclination of the system, which should be explored in future better data. For the low CO model, the associated log and log are 20.5 cm-2 and -2.0 au2, but 17.5 cm-2 and -2.5 au2 for the high model, illustrating the sensitivity to the gas temperature. We note that the CO lines are optically thick in both cases, and that the slab model we used does include the optical depth effects, i.e., the derived and are reliable for the assumed temperature.
@@
We can also estimate the minimum gas mass by assuming the emission is optically thin. Taking the slab model parameters at = 0 km s-1, the total gas mass for CO2, H2O, and NH3 is: 3.21, 6.05 and 5.04 , respectively. Note that the mass for H2O and NH3 should be considered as upper limits given the tentative detection. The total mass for CO is 5.54 for the low case, but 1.75 for the low case. The total gas mass ranges from 4.5 to 5.6 (spreading two orders of mag, depending on the CO mass). Using the Equ. (14) from Birnstiel et al. (2010), we can estimate the Stokes number between the gas and small (0.5 µm) dust by spreading the masses within 0.3 au for simplicity. The Stokes number between the gas and small dust is 0.44 (for the low CO model) or 55 (for the high CO model) assuming a grain density of 3 g cm-3). In the former case, the small dust and gas are coupled and the presence of gas would have an impact on dust dynamics, affecting the inward or outward draft rate depending on the gas surface density. In the later case, the tiny amount of gas has no strong effects on dust, i.e., radiation blowout operates on orbital timescales (less than a year in sub-au).
@@ C.2 Upper Limits on Molecular Hydrogen Gas
No H2 lines are detected in the continuum-subtracted spectrum of HD 23514. We used the local rms estimated around 50 spectral resolution elements around the potential lines to estimate the 3- upper limits on the integrated line flux assuming the lines are optically thin and unresolved. Fixing the gas temperature at 9 00 K, we used all the upper limit line fluxes from H2(S1) to H2(S8) to estimate the gas mass and found that H2(S3) line at 9.664 µm (integrated line flux of 3.27 W cm-2) gives the best 3- upper limit mass, 2.25. The best estimated CO mass is 5.54 (taking from the low 900 K case), implying the CO-to-H2 mass ratio is which is slightly larger than the typical ISM value (an abundance of 10-4 translates to a mass ratio of 1.4. Although the CO-to-H2 mass ratio cannot unambiguously prove the gas is secondary, the old age of the system and the lack of HI and [Ne II] emission favor non-primordial.
@@ C.3 Photodissociation Rate and Molecular Shielding Efficiency
We derived the photodissociation lifetimes for CO, CO2, H2O, and NH3 molecules as a function of stellocentric distance assuming that these molecules are exposed to ultraviolet (UV) photons from both the interstellar radiation field (ISRF) and the star without shielding. The Kurucz model described in Appendix A.1 and the standard ISRF (Draine, 1978; van Dishoeck & Black, 1982) were adopted, and the relevant photodissociation cross sections were taken from the online database of the Leiden Observatory444 https://home.strw.leidenuniv.nl/ẽwine/photo/data/photo/ (Visser et al., 2009; Heays et al., 2017). The result was shown in Figure 5. For all molecules, the UV radiation from the star dominates the photodissociation processes within 10 au. We note that HD 23514 is located in the outer skirt of the Pleiades cluster (far away from the central B-type stars), where the ISRF may be different from that of typical ISM, it does not significantly affect the results within 10 au unless it is very different from the assumed one.
@@
For CO, CO2, H2O, and NH3, UV photons with wavelengths shorter than 109, 115, 180, and 207 nm, respectively, are involved in the photodissociation of the molecules. Unfortunately, there are no UV observations available for HD 23514 that would assess the reliability of the adopted Kurucz model at these wavelengths. Since it is a relatively young F5 type star, it is suspected that it may show magnetic activity and thus probably exhibits excess emission in the UV regime. In the case of our Sun, the Kurucz model underestimates the observed emission increasingly towards the shortest wavelengths555see the measured spectrum from https://home.strw.leidenuniv.nl/ẽwine/photo/data/photo_data/radiation_fields/solar.dat. Therefore, our photodissociation lifetimes derived for HD 23514 should be considered as upper limits.
@@
UV shielding could occur by the molecule itself (self-shielding), by other molecules and atoms, and by the presence of dust. We investigated the shielding factor (efficiency) for CO2 by various gases as a function of the column density of the given gases as shown in Figure 8. Two different radiation fields (ISRF and solar) were used along with the molecular properties from the Leiden Observatory database (https://home.strw.leidenuniv.nl/ẽwine/photo/, Heays et al. 2017). Shielding from CO2 itself is really efficient once the column density reaching 1018cm-2 and can increase its photodissociation lifetime more than tenfold. The dust component of typical debris disks is so dilute that it cannot contribute significantly to the shielding of the molecules. However, in the case of HD 23514, it is conceivable that some parts of the dust disk could be optically thick, which could slightly increase the photodissociation lifetime of the molecules. Nevertheless, the photodissociation lifetime of CO2 with self-shielding is still very short, on the order of a few days at 0.05 au. Assuming the gas remains stable over 15 years between the Spitzer and JWST observation, the shielding from tiny dust needs to be thousandfold more effective than molecular gas shielding.

### v2

@@ A.2 Dust Emissivity and Silica Properties
To further validate our approach, we also generated dust emissivity functions for a young star called LkCa 15. The dust emissivity in LkCa 15 was classified as an ISM-like profile by Watson et al. (2009) where the emission is dominated by pristine silicate dust grains (i.e., amorphous and sub-micron). Therefore, we used its mid-IR archival spectra from Spitzer/IRS and JWST/MRS as the input and constructed the pseudocontinuum using the polynomial function for the former but a combination of two blackbodies for the latter. The resultant two dust emissivity functions agree well within a few percent in the 10 µm complex (5–14 µm), but in the 20 µm complex the JWST one is slightly higher by 20–40% likely due to the limited long-wavelength in determining the pseudocontinuum. The derived JWST LkCa 15 dust emissivity is shown in the right panel of Figure 1. Because our focus is in deriving the crystalline mass fraction indices by referencing the ISM-like profile in the 10 µm complex, the difference in the 20 µm region has no impact on the result. We used the JWST LkCa 15 emissivity function to compute the indices ( for pyroxene, for olivine and for silica) as defined by Watson et al. (2009). Note that these indices are defined as the normalized ratio in dust emissivity at a certain wavelength range covering the narrow features, the normalization of the reference (ISM-like) emissivity (i.e., the blue curve in the left panel of Figure 1) has no effects on the results. For HD 23514, the derived indices are =2.54, =0.91, and 1.8, suggesting a high fraction of crystallinity preferentially in the pyroxene and silica mineralogy. We also derived the dust emissivity using the Spitzer IRS spectrum, and found similar indices within the uncertainty (corroborating no change in the dust composition).
@@
To quantitatively characterize the silica feature in HD 23514, we also derived the dust emissivity using the Spitzer IRS spectrum of T51 (a few Myr-old star). T51 is one of the five silica-rich PPDs characterized by Sargent et al. (2009) who found that the system lacks submicron amorphous dust and the crystalline silicate mass fraction is 15% including both pyroxene and olivine types, and the silica dust is best scribed as annealed silica although obsidian or amorphous quartz give equally good fits. Our derived indices for T51 are =2.02, =1.50, and =2.36, respectively, consistent with the finding for HD 23514 above. Although the 10 µm feature in T51 and HD 23514 peaks at 9.1 µm (Figure 1 right), the FWHM of the 10 µm feature in T51 is 1.5 times larger, suggesting the dominant size in HD 23514 is smaller, likely submicron.
@@
Silica is not abundant in the ISM. Its presence in circumstellar disks requires special conditions for formation like giant impacts as we discuss here. Space weathering, as discussed by Sargent et al. (2009), is one way to deposit silica by converting silicate minerals; however, it is unlikely to be the primary mechanism responsible for the large amount of sub-um size silica seen in HD 23514. Space weathering is a long-term process. µm-size silicate grains in the inner region of HD 23514 have short lifetimes (blowout size is 1 µm, and will be gone on orbital timescales), implying that converting small silicates, produced by collisional cascades, to silica is an inefficient process. For stars just emerge from the protoplanetary disk stage, massive flare events could speed up the process as suggested by Lisse et al. (2020), however, HD 23514 is not very active at an old age of 150 Myr (Appendix A.1). Comets could bring a large amount of fresh dust into the inner region during a period of high delivery rates like the late heavy bombardment, and create silica through space weathering. For solar system comets, the dust mass is 1014 g on average (Harker et al., 2023), meaning a total of 10 billion comets is needed to account for the dust observed in HD 23514 (1024 g derived in Appendix D). The most challenging problem in the comet delivery scenario is that the typical dust size is large, 100 µm (Rinaldi et al., 2017), i.e., space weathering should produce large-size silica, not sub-. In summary, the proposed giant-impact scenario is the most likely explanation for the large amount of submicron silica grains in HD 23514.
Appendix B Other Mid-IR Data and 3–5 µm Photometric Monitoring
@@
HD 23514 has various space- and ground-based mid-IR photometry, with the earliest measurement from IRAS in 1983 (Moshir et al., 1992; Ishihara et al., 2010; Rhee et al., 2008; Wright et al., 2010; Meng et al., 2012). Figure 7 summarizes these photometric observations between 5 and 25 µm. The plotted data have been color corrected using the Spitzer/IRS spectrum, assuming that it is a good representation of the SED shape in the given wavelength range over the period of interest (1983–2010). Because WISE 12 µm band was selected to match approximately that of IRAS (Cutri et al., 2012), they trace the same part of the 10 µm feature. The color corrected flux densities in these WISE and IRAS bands are 102.94.8 mJy and 131.618.4 mJy, respectively, i.e. the two fluxes agree within the 1.5 photometric uncertainties.
@@
IR variability within 20–30% levels, as an additional hallmark commonly found in the extreme debris disks (Moór et al., 2021), is known for HD 23514 (Meng et al., 2012). HD 23514 has been intensively monitoring by Spitzer/IRAC and WISE at 3–5 µm with the former in 3–day cadence when available, while the latter in 6–month cadence. Following the analysis method used for another IR variable, the RZ Psc disk (Su et al., 2023), Figures 8 and 9 show the short- and long-term disk fluxes at 3–5 µm between 2008 and 2024 where the two epochs of the IR spectra are also marked.
@@
Using these 3–5 µm data, the HD 23514 disk exhibits stochastic, both weekly and yearly, IR variability with a weak, positive trend between the excess flux and color temperature. The inferred color temperatures are 700–800 K, consistent with the warmer dust temperature inferred from the featureless pseudocontinuum in Appendix A.2. The typical peak-to-peak flux variation (Figure 8) is 15% in both the 3.6 and 4.5 µm bands, more than 10 significance compared to the photometry stability (1%) determined by field stars in the data. The 14-year long WISE data (Figure 9) show a similar behavior with the largest peak-to-peak flux difference of 25 mJy in the W2 band. Taking the typical (10 mJy) and maximum (25 mJy) 4.5 µm flux change at a face value, the minimal corresponding area change in dust cross section is (1–2.5) au2 by adopting a dust temperature of 800 K and a distance of 139 pc under optically thin assumption. It is interesting to note that this minimal area change in the dust cross section is similar to the hot CO2 gas emitting area (Table 1). The cross section changes can also be translated to the minimal change in the dust mass (assuming static state collisional cascades down to the radiation blowout size of 1 µm). The very lower limit on the associated dust mass is on the order of 3–7 g, corresponding to the total disruption of a 120–165 km-size asteroid (assuming a density of 3.5 g cm-3). This mass/size is a lower limit because of our assumptions (optically thin and static state).
Appendix C Gas Model and Other Related Calculations
C.1 Local Thermodynamic Equilibrium Slab Model
@@
To fit the line emission, we adopted the slab-modeling Python package iris (Munoz-Romero et al., 2023) assuming the gas is in LTE, with line data from the HITRAN database (Gordon et al., 2022). The line width was estimated with thermal broadening based on the temperature of the gas. In iris, the fitting is performed with the Bayesian nested sampling Python package dynesty (Speagle, 2020). We set the stopping criterion by the change in the remaining evidence (marginal likelihood ) when , where = with as the number of parameters in the model.
We restricted the fitting within the wavelength range of 14.0–17.0 µm for the emission of CO2, and adopted uniform priors for emitting area () and column density (), with and . The prior for the gas temperature () was set to be uniform initially, and later set as a normal distribution centered around 500 K with a standard deviation of around 200 K to speed up the search. We included 13CO2 with CO2 in our fitting, and assumed the emissions were from the same region ( and are the same) and the 12C/13C ratio was the same as ISM value (, e.g., Milam et al., 2005). The posterior distributions for the CO2 fitting are shown in Figure 10. The 15.4 µm 13CO2 line is the only (i.e., strongest) isotopologue line present in the data. Excluding it produces a slightly poor fit at 15.4 µm as shown in Figure 2. The best fit CO2 model parameters are summarized in Table 1 by assuming two different turbulence velocities: 0 km s-1 as the nominal and 2 km s-1 as the upper limit value. The two cases make very little difference in the resulting fits visually, illustrating the associated uncertainties in the MCMC fits might be underestimated, and reflecting the limitations of using the simple LTE slab model.
@@
Although there are hints of other molecules (CO, H2O and NH3) present in the continuum-subtracted spectra (Figure 12), we did not attempt to determine their precise parameters using the slab models due to high noise and weak lines. Instead, we assumed they share similar model parameters as CO2, and adjust (and , if necessary) to achieve reasonable fits visually. The mid-IR emission from molecules is due to ro-vibrational or vibrational transitions and is sensitive to their temperatures, i.e., the lines get significantly weaker as the temperature gets cooler. It is then meaningless to explore lower temperatures (or other parameters) without high S/N data. Keeping the same temperature for both H2O and NH3, the column density needs to be reduced by a factor of 2.2 and 1.4, respectively, in order to produce similar flux levels of the lines between 6.5 and 9.5 µm. The low abundances in both H2O and NH3 are consistent with their very short photodissociation lifetimes (Figure 5), by 2–3 order of magnitudes compared to that of CO2, without shielding. The model parameters for CO, which has the longest photodissociation lifetimes near HD 23514, are very uncertain due to the limited wavelength coverage from MIRI/MRS. The slab model has great difficulty in reproducing the observed line width if CO has a similar as CO2 (termed low case in the top panel of Figure 11). Alternatively, the broad width of CO lines can roughly be reproduced by an extra broadening of 200 km s-1 and increasing to 3000 K (termed high case in the middle panel of Figure 11). Such a high temperature can be reached at a stellocentric distance of 0.015 au around HD 23514 where its Keplerian velocity () is 277 km s-1. This seems to suggest the additional broadening might be related to the Keplerian motion of the CO gas. If the gas is indeed in a disk form, resolving the double-peak profiles in the CO emission could help to constrain the inclination of the system, which should be explored in better future data. For the low CO model, the associated log and log are, respectively, 20.5 cm-2 and -2.0 au2, but 17.5 cm-2 and -2.5 au2 for the high model, illustrating the sensitivity to the gas temperature. We note that the CO lines are optically thick in both cases, and that the slab model we used does include optical depth effects, i.e., the derived and are reliable for the assumed temperature.
@@
We can also estimate the minimum gas mass by assuming the emission is optically thin. Taking the slab model parameters at = 0 km s-1, the total gas mass for CO2, H2O, and NH3 is 3.21, 6.05 and 5.04 , respectively. Note that the masses for H2O and NH3 should be considered as upper limits given the tentative detection. The total mass for CO is 5.54 for the low case, but 1.75 for the high case. The total gas mass ranges from 4.5 to 5.6 (spreading two orders of mag, depending on the CO mass). Using the Equ. (14) from Birnstiel et al. (2010), we can estimate the Stokes number between the gas and small (0.5 µm) dust by spreading the masses within 0.3 au for simplicity. The Stokes number between the gas and small dust is 0.44 (for the low CO model) or 55 (for the high CO model) assuming a grain density of 3 g cm-3. In the former case, the small dust and gas are coupled and the presence of gas would have an impact on dust dynamics, affecting the inward or outward draft rates depending on the gas surface density. In the later case, the tiny amount of gas has no strong effects on dust, i.e., radiation blowout operates on orbital timescales (less than a year in the sub-au region).
@@ C.2 Upper Limits on Molecular Hydrogen Gas
No H2 lines are detected in the continuum-subtracted spectrum of HD 23514. We used the local rms estimated around 50 spectral resolution elements around the potential lines to estimate the 3 upper limits on the integrated line flux assuming the lines are optically thin and unresolved. Fixing the gas temperature at 9 00 K, we used all the upper limit line fluxes from H2(S1) to H2(S8) to estimate the gas mass and found that H2(S3) line at 9.664 µm (integrated line flux of 3.27 W cm-2) gives the best 3- upper limit mass, 2.25. The best estimated CO mass is 5.54 (taking from the low 900 K case), implying the CO-to-H2 mass ratio is which is slightly larger than the typical ISM value (an abundance of 10-4 translates to a mass ratio of 1.4. Although the CO-to-H2 mass ratio cannot unambiguously prove the gas is secondary, the old age of the system and the lack of HI and [Ne II] emission favor nonprimordial.
@@ C.3 Photodissociation Rate and Molecular Shielding Efficiency
We derived the photodissociation lifetimes for CO, CO2, H2O, and NH3 molecules as a function of stellocentric distance assuming that these molecules are exposed to ultraviolet (UV) photons from both the interstellar radiation field (ISRF) and the star without shielding. The Kurucz model described in Appendix A.1 and the standard ISRF (Draine, 1978; van Dishoeck & Black, 1982) were adopted, and the relevant photodissociation cross sections were taken from the online database of the Leiden Observatory444 https://home.strw.leidenuniv.nl/ẽwine/photo/data/photo_data/ (Visser et al., 2009; Heays et al., 2017). The result is shown in Figure 5. For all molecules, the UV radiation from the star dominates the photodissociation processes within 10 au. We note that HD 23514 is located in the outskirts of the Pleiades cluster (far away from the central B-type stars), where the ISRF may be different from that of typical ISM, it does not significantly affect the results within 10 au unless its local ISRF is very different from the assumed one.
@@
For CO, CO2, H2O, and NH3, UV photons with wavelengths shorter than 109, 115, 180, and 207 nm, respectively, are involved in the photodissociation of the molecules. Unfortunately, there are no UV observations available for HD 23514 that would assess the reliability of the adopted Kurucz model at these wavelengths. Since it is a relatively young F5-type star, it is suspected that it may show magnetic activity and thus probably exhibits excess emission in the UV regime. In the case of our Sun, the Kurucz model underestimates the observed emission increasingly toward the shortest wavelengths555see the measured spectrum from https://home.strw.leidenuniv.nl/ẽwine/photo/data/photo_data/radiation_fields/solar.dat. Therefore, our photodissociation lifetimes derived for HD 23514 should be considered as upper limits.
@@
UV shielding could occur by the molecule itself (self-shielding), by other molecules and atoms, and by the presence of dust. We investigated the shielding factor (efficiency) for CO2 by various gases as a function of the column density of the given gases as shown in Figure 8. Two different radiation fields (ISRF and solar) were used along with the molecular properties from the Leiden Observatory database (https://home.strw.leidenuniv.nl/ẽwine/photo/, Heays et al. 2017). Shielding from CO2 itself is really efficient once the column density reaches 1018cm-2, which can increase its photodissociation lifetime by more than tenfold. The dust component of typical debris disks is so dilute that it cannot contribute significantly to the shielding of the molecules. However, in the case of HD 23514, it is conceivable that some parts of the dust disk could be optically thick, which could slightly increase the photodissociation lifetime of the molecules. Nevertheless, the photodissociation lifetime of CO2 with self-shielding is still very short, on the order of a few days at 0.05 au. Assuming the gas remains stable over 15 yr between the Spitzer and JWST observation, the shielding from tiny dust needs to be thousandfold more effective than molecular gas shielding.